OpenStudy (anonymous):

Find all integer solutions to 14x + 77y = 70

OpenStudy (cwrw238):

7(2x + 11y) = 70 2x + 11y = 10 lol - i'm lost now if there are solutions then 1 must be negative

jimthompson5910 (jim_thompson5910):

It's better to solve 2x + 11y = 1 first, then scale both sides appropriately.

jimthompson5910 (jim_thompson5910):

Notice how 11 = 2*5 + 1, so 11 = 2*5 + 1 11*1 = 2*5 + 1 11*1 - 2*5 = 1 2(-5) + 11(1) = 1 This means that the solutions for 2x+11y = 1 are x = -5 and y = 1

jimthompson5910 (jim_thompson5910):

Once you know that 2(-5) + 11(1) = 1 multiply both sides by 10 2(-5) + 11(1) = 1 10( 2(-5) + 11(1)) = 1*10 10*2*(-5) + 10*11*(1) = 10 2*(-5*10) + 11*(10*1) = 10 2*(-50) + 11*(10) = 10 So the solutions to 2x+11y = 10 are x = -50 and y = 10

jimthompson5910 (jim_thompson5910):

That's just one integer solution, but you can use it to find the rest.

OpenStudy (maheshmeghwal9):

@jim_thompson5910 Why did u nt take 10 at RHS. Plz tell:)

jimthompson5910 (jim_thompson5910):

because it's much easier to solve the equation when the RHS is 1, then you scale appropriately

OpenStudy (maheshmeghwal9):

oh i see. thnx:)

jimthompson5910 (jim_thompson5910):

It's shown in the work above

OpenStudy (cwrw238):

clever stuff!

OpenStudy (maheshmeghwal9):

of course:) Previous was much clever:)

OpenStudy (asnaseer):

\[\begin{align} 14x+77y&=70\\ \therefore2x+11y&=10\\ \therefore11y&=10-2x=2(5-x)\\ \therefore y&=\frac{2(5-x)}{11}\implies5-x=11n\text{ (i.e. (5-x) must be a multiple of 11)}\\ \therefore y&=\frac{2*11n}{11}=2n \end{align}\]so solutions are:\[x=5-11n\]and\[y=2n\]for:\[n=-\infty\text{ to }+\infty\]

OpenStudy (anonymous):

Here's something that our tutor thought was the right answer. x=-50+11k y=10-2k for all k integers

OpenStudy (asnaseer):

there are many forms for the /correct/ answer. they all generate the same pair of x and y values

OpenStudy (asnaseer):

e.g. this is what wolfram comes up with: http://www.wolframalpha.com/input/?i=solve+14x+%2B+77y+%3D+70+over+integers

OpenStudy (asnaseer):

wolfram's solution uses an 'n' that is the negative of my 'n'

OpenStudy (anonymous):

ok, back to the other question...with a negative sign, are you familiar with the continued fraction method?

OpenStudy (asnaseer):

lets go back there...